# Formulation of entropy-conservative discretizations for compressible flows of thermally perfect gases

### v2

@@ 1 Introduction
To address these challenges, research has been devoted to devising numerical discretizations capable of automatically reproducing important features of the continuous equations, such as enforcing physical symmetries or conserving primary or secondary invariants of the considered equations [2]. These schemes, usually referred to as structure-preserving methods, have been shown to provide increased robustness and reliability, and have been the subject of many studies in recent years ([3, 4, 5, 6, 7, 8, 9])
@@
In incompressible flows, kinetic energy acts as a bounded $L_{2}$ measure of velocity [10], and as such it acts as a mathematical entropy for the incompressible Euler equations and can be used to assess nonlinear stability. In the context of compressible flow simulations, kinetic energy is no longer a conserved quantity and cannot provide a bound on the solution, but still the use of Kinetic Energy Preserving (KEP) methods has been successfully adapted from the incompressible case. These methods are the most established and widely used among structure-preserving schemes [11, 12, 8]. The conditions to obtain KEP schemes have been defined both in the context of finite volume methods [13, 14] and finite difference methods [8, 15]. However, the most commonly used definition for the KEP property introduces some indeterminacy, due to the arbitrariness of the discretization of the pressure term  [16, 17], which nevertheless has been found to have a significant impact on the evolution of kinetic energy [18].
@@
In compressible flows, kinetic energy no longer serves as a mathematical entropy of the governing equations; instead, in the inviscid limit, this role is primarily assumed by the thermodynamic entropy, whose existence is closely linked to the symmetrization of the Euler equations [19, 20]. Accordingly, another important class of methods is that based on Entropy Conservative (EC) discretizations [21, 22, 23], which guarantee a correct induced discrete balance of entropy in absence of discontinuities. A fundamental difference between the two formulations is that, unlike KEP schemes, which only depend on the discretization of the continuity and momentum equations, EC schemes depend on the discretization of the energy equation, which, in turn, depends on the Equation of State (EoS) assumed for the gas.
@@
In the case of a calorically perfect gas, several methods have been developed that can preserve the entropy balance at the discrete level both exactly ([21, 24, 23, 25]) or approximately with arbitrary accuracy ([26, 27, 28]). Most EC schemes have been developed within the framework of the technique introduced by Tadmor [29], which has been applied to finite volume and finite element schemes. An improvement in the affordability of EC schemes came from the work of Ismail and Roe [22], which introduced the use of fluxes based on the logarithmic mean serving as the foundation of many subsequent developments. Building on this, Chandrashekar [24] proposed the first scheme that simultaneously satisfied both the KEP and the EC property. More recently, Ranocha [23, 30, 16] introduced a different KEP and EC formulation which, among other advantages, was shown to have a discretization of the pressure which led to more physically consistent treatment of kinetic energy [18].
@@
An alternative gas model is that of the thermally perfect gas, in which it is taken into account the temperature variation of specific heats caused by the excitation of vibrational energy within the molecules of the gas and by the electronic energy associated with electron motion within the atoms and molecules [31]. The use of this model more accurately describes the behavior at the high temperatures typical of combustion phenomena [32] as well as many wall-bounded high-speed flows, when temperature may exhibit both high absolute values and wide ranges [33, 34, 35]. In recent years, some extensions of the EC schemes to the case of thermally perfect gases have also been developed [36, 37, 38]. Notably, all these formulations reduce to the scheme of Chandrashekar [24] in the limit of a calorically perfect gas, owing to an analogous treatment of the pressure terms.
@@
The structure of the paper is as follows. In Section 2, we introduce the governing equations, including the thermodynamic closure for thermally perfect gases. We then recall the necessary conditions for kinetic energy preservation and entropy conservation, and show how the general EC condition particularizes to this case. A polynomial-based model for the temperature dependence of specific heats is adopted, leading to a discrete formulation that is EC, KEP, and free of singularities. A comparison with existing schemes is also presented.
@@ 2 Problem formulation
The analysis is conducted within the framework of a conservative semidiscretized treatment of the governing equations, in which the time derivatives of the spatially discretized variables are driven by the difference of numerical fluxes at adjacent faces, which comprise convective and pressure contributions consistent with the components of $\mathbf{f}$ in Eq. (1). We work in a Finite Difference (FD) framework, although the numerical fluxes determined in Sec. 2.1 and 3.1 can also be used with other discretization techniques (e.g. with Finite Volume or Discontinuous Galerkin formulations). The method will be illustrated for second-order two-point fluxes on a uniform grid $x_{i}$ of width $h=x_{i+1}-x_{i}$; the corresponding high-order extension can be constructed through the approach used in [26, 39, 40]. The numerical flux for a generic quantity $\rho\phi$ at the interface between cell $i$ and $i+1$ will be indicated by $\mathcal{F}_{\rho\phi}$, whereas its convective contribution will be denoted by $\mathcal{F}^{\text{c}}_{\rho\phi}$. All the schemes we will consider in the analysis are KEP, which implies that the convective term in the momentum flux is given by the product between the mass flux $\mathcal{F}_{\rho}$ and the arithmetic average of velocity $\overline{u}=(u_{i}+u_{i+1})/2$. This is indeed the necessary and sufficient condition for (second-order, two-point) fluxes to be KEP [13, 14]. An important assumption we make in this study is that the pressure term in the momentum equation is discretized using a second-order central derivative, giving rise to the general form for the momentum flux: $\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\overline{u}\,+\overline{p}\,$. As it is known, this formulation induces a locally conservative discretization of the convective term in the kinetic-energy equation, with flux $\mathcal{F}^{\text{c}}_{\rho k}=\mathcal{F}_{\rho}u_{i}u_{i+1}/2$  [15], and a central approximation of the non conservative pressure term in the form $u\partial p/\partial x\rightarrow u\delta p/\delta x$ with $\delta p=p_{i+1}-p_{i-1}$ and $\delta x=x_{i+1}-x_{i-1}=2h$.
@@ 2.2 Polynomial-based fitting model
In Eq. (7) $\ell$ and $r$ are arbitrary natural numbers, whose values are used to give flexibility to the fitting capability of the functional dependence $c_{v}(T)$. We will detail the derivation of the EC fluxes for this general formulation, even though usually only 5, 7 or 9 coefficients are used to experimentally fit the gas behavior [42, 43]. In addition to being widely used, this form of $c_{v}(T)$ is also easily tractable from an analytical point of view. For another example of functional dependence $c_{v}(T)$, see Sec. 3.1.
@@
$\dfrac{\delta^{\,+}\omega/T}{\delta^{\,+}1/T}=\varepsilon_{\mathrm{ref}}+c_{-1}\left(1+\dfrac{\delta^{\,+}\frac{\log{T}}{T}}{\delta^{\,+}1/T}\right)+c_{0}\dfrac{\delta^{\,+}\log{1/T}}{\delta^{\,+}1/T}-\sum_{\begin{subarray}{c}m=-\ell\\ m\neq-1,0\end{subarray}}^{r}\frac{c_{m}}{m(m+1)}\frac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T}.$ (8)
in which
$\varepsilon_{\mathrm{ref}}=e_{\mathrm{ref}}-c_{-1}\log{T_{\mathrm{ref}}}-\sum_{\begin{subarray}{c}m=-\ell\\ m\neq-1\end{subarray}}^{r}\frac{c_{m}}{m+1}T_{\mathrm{ref}}^{m+1}$
and it is equivalent to the $b_{1}$ coefficient present in NASA polynomials in [43]. The variable term associated with $c_{-1}$ in Eq. (8) can be manipulated as
@@
The scheme based on Eqs. (2), (6) and (10) is exactly entropy conserving. The use of the mass flux $\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,$ has reduced, in the case of thermally perfect gases, the singularity of the original formulation for an arbitrary EoS to that associated with the computation of the logarithmic mean, for which a local fix analogous to that devised by Ismail and Roe [22] can be used. Asymptotically Entropy Conserving (AEC) formulations similar to that developed by De Michele and Coppola [26] could also be easily derived by using the asymptotic expansion of the logarithm employed in the calorically perfect case. In fact, using the Taylor series expansion
@@ 2.3 Comparison with existing schemes
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}},\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\varepsilon_{\mathrm{ref}}+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{r}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}+\dfrac{u_{i}u_{i+1}}{2}\right]+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overline{u}\,\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}}.$ (14)
@@
By comparing Eqs. (13) and (14) one observes that the two sets of fluxes are almost coincident, except for the pressure terms in the momentum and total-energy equations, highlighted in red in Eq. (14). While in the present formulation (Eq. (13)) the pressure flux in the momentum equation is expressed through the arithmetic mean $\overline{p}\,$, in the formulation by Gouasmi et al. (Eq. (14)) it is expressed by using the complex average $\widehat{p}=R\,\overline{\rho}\,/\overline{1/T}\,$, which depends on density and temperature. Moreover, the pressure term in the total-energy equation is expressed as $\overline{\overline{(p,u)}}\,$ in Eq. (13) and as $\widehat{p}\,\,\overline{u}\,$ in Eq. (14). It is well known that the discretization of the pressure term in the momentum equation plays an important role in the global discrete preservation of kinetic energy. In fact, in [18], which deals with KEP split forms for calorically perfect gases, it is shown that even in the case in which the convective flux in the momentum equation is written in the KEP form as $\mathcal{F}^{\text{c}}_{\rho u}=\mathcal{F}_{\rho}\overline{u}\,$, the discretization of the pressure term can spoil the correct balance of induced kinetic energy causing errors in the pressure work contribution. This is especially true when, as in the case of the Chandrashekar scheme [24] to which Eq. (14) reduces for ideal gases, the pressure average is rendered as a combination of averages of density and temperature. The conclusions of the analysis conducted in [18] are that, in the case of calorically perfect gases, the simple arithmetic average of pressure in the momentum flux is the choice yielding the best results for the case of inviscid Taylor–Green vortices. In Sec. 4.2 we show that this result also applies in the case of thermally perfect gases. Similarly, the treatment of the pressure term in the total energy equation is known to be an element potentially affecting certain statistical quantities related to the fluctuating motion in homogeneous isotropic turbulence. In [17] and [12], as an example, it is shown that r.m.s of density and temperature fluctuations in homogeneous isotropic turbulence simulations can erroneously grow without bound if the pressure term in the total energy equation is different from $\overline{\overline{(p,u)}}\,$. In Sec. 4.2 we show that this is the case also for thermally perfect gases, demonstrating the importance of the set of fluxes here derived as a relevant improvement on existing formulations.
@@ 3.1 Numerical fluxes for hypersonic flows with the RRHO model
The general methodology illustrated in Sec. 2.1 can be used to derive EC numerical fluxes for arbitrary thermally perfect models in which the constant-volume specific heat $c_{v}(T)$ is amenable to an analytical treatment. In this section, we give a further example of the procedure by analyzing the Rigid-Rotor Harmonic-Oscillator (RRHO) model used for hypersonic flows’ simulations as reported by [37] and adapted from a formulation presented in [44]. The model takes into account vibrational and rotational contributions to internal energy, which for a single diatomic gas takes the form
@@ 3.2 Multicomponent flows
In the case of a gas composed of $N_{s}$ nonreacting species, the usual standard approach considers the mixture as a single compressible fluid, with density, momentum and total energy (per unit volume) given by the sum of the individual components, where the density of each species evolves according to a separate continuity equation. In this case, the compressible Euler equations are still in the form of Eq. (1) where $\mathbf{u}=\left(\rho_{1},\rho_{2},\ldots,\rho_{N},\rho u,\rho E\right)^{\text{T}}$ and $\mathbf{f}=\left(\rho_{1}u,\rho_{2}u,\ldots,\rho_{N_{s}}u,\rho uu+p,u(\rho E+p)\right)^{\text{T}}$, where $\rho_{k}=\rho Y_{k}$ is the density of the $k$-th species, $Y_{k}$ being the mass fraction and $\rho$ the density of the mixture. From the relation $\sum_{k}Y_{k}=1$ it follows $\sum_{k}\rho_{k}=\sum_{k}\rho Y_{k}=\rho$. Internal energy and entropy of the mixture are defined as $\rho e=\sum_{k}\rho_{k}e_{k}$ and $\rho s=\sum_{k}\rho_{k}s_{k}$, which also implies $e=\sum_{k}Y_{k}e_{k}$ and $s=\sum_{k}Y_{k}s_{k}$. Individual equations of state $p_{k}=\rho_{k}R_{k}T$, where $R_{k}$ is the gas constant of the $k$-th species, and Dalton’s law $\sum_{k}p_{k}=p$ are assumed. The Gibbs’ relation, which holds for each species: $d\left(\rho_{k}e_{k}\right)=Td\left(\rho_{k}s_{k}\right)+g_{k}d\rho_{k}$, can be used to relate the convective and pressure terms in the balance equations for mass, internal energy and entropy for each component of the mixture, as it was done for a single compound in [39]. By imposing a locally conservative formulation for mass and internal energy, and a straightforward discretization of the pressure term, one obtains the analogous sufficient conditions for entropy conservation of each species:
@@
which is the extension of the Ranocha’s flux to multicomponent flows of ideal gases. To the best of the authors’ knowledge, this extension has not been mentioned in the open literature up to now.

### v3

@@ 1 Introduction
To address these challenges, research has been devoted to devising numerical discretizations capable of automatically reproducing important features of the continuous equations, such as enforcing physical symmetries or conserving primary or secondary invariants of the considered equations [2], or by imposing invariant domain constraints, such as positivity of the density or internal energy [3, 4]. These schemes, usually referred to as structure-preserving methods, have been shown to provide increased robustness and reliability, and have been the subject of many studies in recent years [5, 6, 7, 8, 9].
@@
In incompressible flows, kinetic energy acts as a bounded $L_{2}$ measure of velocity [10], and as such it acts as a mathematical entropy for the incompressible Euler equations and can be used to assess nonlinear stability. In the context of compressible flow simulations, kinetic energy is no longer a conserved quantity and cannot provide a bound on the solution, but still the use of Kinetic Energy Preserving (KEP) methods has been successfully adapted from the incompressible case. These methods are the most established and widely used among structure-preserving schemes [11, 12, 9, 13]. The conditions to obtain KEP schemes have been defined both in the context of finite volume methods [14, 15] and finite difference methods [9, 16]. However, the most commonly used definition for the KEP property introduces some indeterminacy, due to the arbitrariness of the discretization of the pressure term  [17, 18], which nevertheless has been found to have a significant impact on the evolution of kinetic energy [19].
@@
In compressible flows, kinetic energy is not a mathematical entropy of the governing equations; instead, in the inviscid limit, this role is primarily assumed by the thermodynamic entropy, whose existence is closely linked to the symmetrization of the Euler equations [20, 21]. Accordingly, another important class of methods is that based on Entropy Conservative (EC) discretizations [22, 23, 24], which guarantee a correct induced discrete balance of entropy in absence of discontinuities. These methods have been typically used as baseline formulations to design entropy stable discretizations, which enforce semi-discrete entropy inequality [25]. This approach complements the more classical techniques based on artificial (or entropy) viscosity [26, 27, 28, 29, 30]. A fundamental difference between KEP and EC formulations is that, unlike KEP schemes, which only depend on the discretization of the continuity and momentum equations, EC schemes depend on the discretization of the energy equation, which, in turn, depends on the Equation of State (EoS) assumed for the gas.
@@
In the case of a calorically perfect gas, several methods have been developed that can preserve the entropy balance at the discrete level both exactly [22, 31, 24, 32] or approximately with arbitrary accuracy [33, 34, 35]. Most EC schemes have been developed within the framework of the technique introduced by Tadmor [25], which has been applied to finite volume and finite element schemes. An improvement in the affordability of EC schemes came from the work of Ismail and Roe [23], which introduced the use of fluxes based on the logarithmic mean serving as the foundation of many subsequent developments. Building on this, Chandrashekar [31] proposed the first scheme that simultaneously satisfied both the KEP and the EC property. More recently, Ranocha [24, 36, 17] introduced a different KEP and EC formulation which, among other advantages, was shown to have a discretization of the pressure which led to more physically consistent treatment of kinetic energy [19].
@@
An alternative gas model is that of the thermally perfect gas, in which it is taken into account the temperature variation of specific heats caused by the excitation of vibrational energy within the molecules of the gas and by the electronic energy associated with electron motion within the atoms and molecules [37]. The use of this model more accurately describes the behaviour at the high temperatures typical of combustion phenomena [38] as well as many wall-bounded high-speed flows, when temperature may exhibit both high absolute values and wide ranges [39, 40, 41]. In recent years, some extensions of the EC schemes to the case of thermally perfect gases have also been developed [42, 43, 44]. Notably, all these formulations reduce to the scheme of Chandrashekar [31] in the limit of a calorically perfect gas, owing to an analogous treatment of the pressure terms.
@@
The structure of the paper is as follows. In Section 2, we introduce the governing equations, including the thermodynamic closure for thermally perfect gases. We then recall the necessary conditions for kinetic energy preservation and entropy conservation, and show how the general EC condition particularizes to this case. A polynomial-based model for the temperature dependence of specific heats is adopted, leading to a discrete formulation that is EC, KEP, and essentially free of singularities. A comparison with existing schemes is also presented.
@@ 2 Problem formulation
The analysis is conducted within the framework of a conservative semidiscretized treatment of the governing equations, in which the time derivatives of the spatially discretized variables are driven by the difference of numerical fluxes at adjacent faces, which comprise convective and pressure contributions consistent with the components of $\mathbf{f}$ in Eq. (1). We work in a Finite Difference (FD) framework, although the numerical fluxes determined in Section 2.1 and 3.1 can also be used with other discretization techniques (e.g. with Finite Volume or Discontinuous Galerkin formulations). The method will be illustrated for second-order two-point fluxes on a uniform grid $x_{i}$ of width $h=x_{i+1}-x_{i}$; the corresponding high-order extension can be constructed through the approach used in [33, 45, 46]. The numerical flux for a generic quantity $\rho\phi$ at the interface between cell $i$ and $i+1$ will be indicated by $\mathcal{F}_{\rho\phi}$, whereas its convective contribution will be denoted by $\mathcal{F}^{\text{c}}_{\rho\phi}$. All the schemes we will consider in the analysis are KEP, which implies that the convective term in the momentum flux is given by the product between the mass flux $\mathcal{F}_{\rho}$ and the arithmetic average of velocity $\overline{u}=(u_{i}+u_{i+1})/2$. This is indeed the necessary and sufficient condition for (second-order, two-point) fluxes to be KEP [14, 15]. An important assumption we make in this study is that the pressure term in the momentum equation is discretized using a second-order central derivative, giving rise to the general form for the momentum flux: $\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\overline{u}\,+\overline{p}\,$. As it is known, this formulation induces a locally conservative discretization of the convective term in the kinetic-energy equation, with flux $\mathcal{F}^{\text{c}}_{\rho k}=\mathcal{F}_{\rho}u_{i}u_{i+1}/2$  [16], and a central approximation of the non conservative pressure term in the form $u\partial p/\partial x\rightarrow u\delta p/\delta x$ with $\delta p=p_{i+1}-p_{i-1}$ and $\delta x=x_{i+1}-x_{i-1}=2h$.
@@ 2.2 Polynomial-based fitting model
In Eq. (7) $\ell$ and $r$ are arbitrary natural numbers, whose values are used to give flexibility to the fitting capability of the functional dependence $c_{v}(T)$. We will detail the derivation of the EC fluxes for this general formulation, even though usually only 5, 7 or 9 coefficients are used to experimentally fit the gas behaviour [48, 49]. In addition to being widely used, this form of $c_{v}(T)$ is also easily tractable from an analytical point of view. For another example of functional dependence $c_{v}(T)$, see Section 3.1.
@@
$\dfrac{\delta^{\,+}\omega/T}{\delta^{\,+}1/T}=\varepsilon_{\mathrm{ref}}+c_{-1}\left(1+\dfrac{\delta^{\,+}\frac{\log{T}}{T}}{\delta^{\,+}1/T}\right)+c_{0}\dfrac{\delta^{\,+}\log{1/T}}{\delta^{\,+}1/T}-\sum_{\begin{subarray}{c}m=-\ell\\ m\neq-1,0\end{subarray}}^{r}\frac{c_{m}}{m(m+1)}\frac{\delta^{\,+}T^{m}}{\delta^{\,+}1/T},$ (8)
where
$\varepsilon_{\mathrm{ref}}=e_{\mathrm{ref}}-c_{-1}\log{T_{\mathrm{ref}}}-\sum_{\begin{subarray}{c}m=-\ell\\ m\neq-1\end{subarray}}^{r}\frac{c_{m}}{m+1}T_{\mathrm{ref}}^{m+1},$
which is equivalent to the $b_{1}$ coefficient present in NASA polynomials in [49]. The variable term associated with $c_{-1}$ in Eq. (8) can be manipulated as
@@
The scheme based on Eqs. (2), (6) and (10) is exactly entropy conserving for thermally perfect gases, and will be denoted as EC-TP. The use of the mass flux $\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,$ has reduced, in the case of thermally perfect gases, the singularity of the original formulation for an arbitrary EoS to that associated with the computation of the logarithmic mean, for which a local fix analogous to that devised by Ismail and Roe [23] can be used. Asymptotically Entropy Conserving (AEC) formulations similar to that developed by De Michele and Coppola [33] could also be easily derived by using the asymptotic expansion of the logarithm employed in the calorically perfect case. In fact, using the Taylor series expansion
@@ 2.3 Comparison with existing schemes
$\mathcal{F}_{\rho}=\overline{\rho}^{\text{log}}\overline{u}\,\qquad\mathcal{F}_{\rho u}=\mathcal{F}_{\rho}\,\overline{u}+{\color[rgb]{1,0,0}\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}},\qquad\mathcal{F}_{\rho E}=\mathcal{F}_{\rho}\left[\varepsilon_{\mathrm{ref}}+c_{0}\dfrac{1}{\overline{1/T}^{\text{log}}}+\sum_{\begin{subarray}{c}m=1\end{subarray}}^{r}\frac{c_{m}}{m+1}\overline{\overline{T}}^{m}+\dfrac{u_{i}u_{i+1}}{2}\right]+{\color[rgb]{1,0,0}\overline{u}\,\dfrac{R\,\overline{\rho}\,}{\overline{1/T}\,}}.$ (14)
@@
By comparing Eqs. (13) and (14) one observes that the two sets of fluxes are almost coincident, except for the pressure terms in the momentum and total-energy equations, highlighted in red in Eq. (14). While in the present formulation (Eq. (13)) the pressure flux in the momentum equation is expressed through the arithmetic mean $\overline{p}\,$, in the formulation by Gouasmi et al. (Eq. (14)) it is expressed by using the complex average $\widehat{p}=R\,\overline{\rho}\,/\overline{1/T}\,$, which depends on density and temperature. Moreover, the pressure term in the total-energy equation is expressed as $\overline{\overline{(p,u)}}\,$ in Eq. (13) and as $\widehat{p}\,\,\overline{u}\,$ in Eq. (14). It is well known that the discretization of the pressure term in the momentum equation plays an important role in the global discrete preservation of kinetic energy. In fact, in [19], which deals with KEP split forms for calorically perfect gases, it is shown that even in the case in which the convective flux in the momentum equation is written in the KEP form as $\mathcal{F}^{\text{c}}_{\rho u}=\mathcal{F}_{\rho}\overline{u}\,$, the discretization of the pressure term can spoil the correct balance of induced kinetic energy causing errors in the pressure work contribution. This is especially true when, as in the case of the Chandrashekar scheme [31] to which Eq. (14) reduces for ideal gases, the pressure average is rendered as a combination of averages of density and temperature. The conclusions of the analysis conducted in [19] are that, in the case of calorically perfect gases, the simple arithmetic average of pressure in the momentum flux is the choice yielding the best results for the case of inviscid Taylor–Green vortices. In Section 4.2 we show that this result also applies in the case of thermally perfect gases. Similarly, the treatment of the pressure term in the total energy equation is known to be an element potentially affecting certain statistical quantities related to the fluctuating motion in homogeneous isotropic turbulence. In [18] and [12], as an example, it is shown that r.m.s of density and temperature fluctuations in homogeneous isotropic turbulence simulations can erroneously grow without bound if the pressure term in the total energy equation is different from $\overline{\overline{(p,u)}}\,$. In Section 4.2 we show that this is the case also for thermally perfect gases, demonstrating the importance of the set of fluxes here derived as a relevant improvement on existing formulations.
@@ 3.1 Numerical fluxes for hypersonic flows with the RRHO model
The general methodology illustrated in Section 2.1 can be used to derive EC numerical fluxes for arbitrary thermally perfect models in which the constant-volume specific heat $c_{v}(T)$ is amenable to an analytical treatment. In this section, we give a further example of the procedure by analyzing the Rigid-Rotor Harmonic-Oscillator (RRHO) model used for hypersonic flows’ simulations as reported by [43] and adapted from a formulation presented in [50]. The model takes into account vibrational and rotational contributions to internal energy, which for a single diatomic gas takes the form
@@ 3.2 Multicomponent flows
In the case of a gas composed of $N_{s}$ nonreacting species, the usual standard approach considers the mixture as a single compressible fluid, with density, momentum and total energy (per unit volume) given by the sum of the individual components, where the density of each species evolves according to a separate continuity equation [51, 42, 52, 53]. In this case, the compressible Euler equations are still in the form of Eq. (1) where $\mathbf{u}=\left(\rho_{1},\rho_{2},\ldots,\rho_{N},\rho u,\rho E\right)^{\text{T}}$ and $\mathbf{f}=\left(\rho_{1}u,\rho_{2}u,\ldots,\rho_{N_{s}}u,\rho uu+p,u(\rho E+p)\right)^{\text{T}}$, where $\rho_{k}=\rho Y_{k}$ is the density of the $k$-th species, $Y_{k}$ being the mass fraction and $\rho$ the density of the mixture. From the relation $\sum_{k}Y_{k}=1$ it follows $\sum_{k}\rho_{k}=\sum_{k}\rho Y_{k}=\rho$. Internal energy and entropy of the mixture are defined as $\rho e=\sum_{k}\rho_{k}e_{k}$ and $\rho s=\sum_{k}\rho_{k}s_{k}$, which also implies $e=\sum_{k}Y_{k}e_{k}$ and $s=\sum_{k}Y_{k}s_{k}$. Individual equations of state $p_{k}=\rho_{k}R_{k}T$, where $R_{k}$ is the gas constant of the $k$-th species, and Dalton’s law $\sum_{k}p_{k}=p$ are assumed. The Gibbs’ relation, which holds for each species: $d\left(\rho_{k}e_{k}\right)=Td\left(\rho_{k}s_{k}\right)+g_{k}d\rho_{k}$, can be used to relate the convective and pressure terms in the balance equations for mass, internal energy and entropy for each component of the mixture, as it was done for a single compound in [45]. By imposing a locally conservative formulation for mass and internal energy, and a straightforward discretization of the pressure term, one obtains the analogous sufficient conditions for entropy conservation of each species:
@@
which is a possible extension of the Ranocha’s flux to multicomponent flows of ideal gases.